OpenStudy (anonymous):

The scale of a certain map is 3/4 inch = 12 miles. What is the actual area, in square miles, of a parking lot represented on the map by a square with sides of 5/8 inch

OpenStudy (katrinakaif):

10 Miles Do you want the solution as well?

OpenStudy (anonymous):

yes please

OpenStudy (katrinakaif):

Okay: Follow on the next post

OpenStudy (katrinakaif):

3/4 = .75 = 12 miles As a ratio, .75/12 = 5/8 over x 5/8 = .625 Then cross multiply (.75 * x) and (12 * .625) = .75x = 75 - Divide by .75 = x = 10

OpenStudy (kfujioka):

Katrina's work is correct in finding the actual length of the parking lot represented by that 5/8 inch square. However the question asks for the area of the actual parking lot. So, as to not mislead you, the full answer to your question is 100 sq. miles.

OpenStudy (katrinakaif):

Thanks for finding the Mistake. Unfortunately I don't have access to the 'Good Answer' Bbutton.. :(